OpenStudy (anonymous):

Simplify. -12*1.6*3+48/3

OpenStudy (mathstudent55):

First, do multiplications and divisions in the order they appear from left to right.

OpenStudy (mathstudent55):

\(-12 \times 1.6 \times 3 + 48 \div 3\) \(= \color{red}{-12 \times 1.6} \times 3 + \color{red}{48 \div 3}\) Start with the two operations in red. Copy everything else down tot eh next line.

OpenStudy (mathstudent55):

Wait. I see the pic now.

OpenStudy (mathstudent55):

\(-12 \times \dfrac{1}{6} \times 3 + 48 \div 3\) \(=\color{red}{-12 \times \dfrac{1}{6}} \times 3 + \color{red}{48 \div 3}\) Do the two operations in red. Copy everything else as it is to the next line.

OpenStudy (anonymous):

so 19.2 and 16

OpenStudy (anonymous):

this is for my semester assesment

pooja195 (pooja195):

Want all the answers?

pooja195 (pooja195):

Here are all the answers to your quiz/test! I took the same exact one :D and got 100% my teacher was proud of me :') I'm sure your teacher will be happy to see you got a 100 too! Now hurry and copy this! http://pages.citebite.com/a4o3b6r4a2tjy

OpenStudy (mathstudent55):

-12 * 1/6 = -2. Where did you get the 19.2 from?

pooja195 (pooja195):

Have fun!

OpenStudy (mathstudent55):

\(=\color{red}{-12 \times \dfrac{1}{6}} \times 3 + \color{red}{48 \div 3}\) Now you get this: \(=-2 \times 3 + 16\) Now we do the multiplication: \(= - 6 + 16\) Finally we add: \(= 10\)

pooja195 (pooja195):

@mathstudent55 we dont help on tests.

OpenStudy (mathstudent55):

I had no idea this was a test. I don't see that stated anywhere.